OpenStudy (anonymous):

What is the math called when you are finding the location of the Y and x Axis

OpenStudy (anonymous):

Do you mean cootdinates? The x and y axis is the line of all points within the specified range for x and y

OpenStudy (anonymous):

coordinates*

OpenStudy (anonymous):

Yes I forgot how to work out the problem for finding the Accordance to make the lines "I'm studying for the act"

OpenStudy (anonymous):

Let's work on the vocab and terminology here. If you are using a coordinate system (x- and y-axes) to locate a point, you are using coordinate geometry (sometimes called "analytic geometry" depending on what you're doing).

OpenStudy (anonymous):

Okay I Do you remember shading in a part of the graph depending on which was greater than the other?, Having certain numbers on each side of the = cancel out themselves using division multiplication subtraction, Addition

OpenStudy (anonymous):

It sounds like you're talking about solving inequalities, but it's hard to tell exactly what you mean.

OpenStudy (anonymous):

I wish I could take a picture of what I'm trying to learn in my ACT book but I don't think I can do that. How would you Get the Y and ask Coordinates for 3X + 2Y =21 , Then decide if it's a dotted line or solid line an which have to Shade in? That's what I'm doing

OpenStudy (anonymous):

Ah. Ok. It looks like you have a linear inequality. You only have an equals sign between the 3x+2y and the 21; is that supposed to be an inequality sign of some kind? ≥ or ≤? < or >?

OpenStudy (anonymous):

Yes sir it is

OpenStudy (anonymous):

You'll first want to solve the inequality for y, to get it into y=mx+b form, then the shading is down in the direction of the inequality: up for > and down for <.

OpenStudy (anonymous):

I believe I have forgotten how to solve Y = MX+B

OpenStudy (anonymous):

It's the same as any other kind of solving: You get what you're solving for on one side of the equals sign and move everything else over to the other side. You'll want to review these basics http://themathpage.com/alg/equations.htm And then review linear equations in general http://themathpage.com/alg/equation-of-a-line.htm